OpenStudy (anonymous):

help!! WILL MEDAL What is the slope-intercept form of the function that contains the point (1, –2) and has a slope of –3? y =________ x + ________

OpenStudy (trojanpoem):

y+2/ x - 1 = -3 can you continue ?

OpenStudy (anonymous):

i'm not sure

OpenStudy (trojanpoem):

y+2 = -3x +3 y = -3 x + 1

OpenStudy (anonymous):

sorry, i still don't get it

OpenStudy (trojanpoem):

The slope of a line is y- y(1) / x- x(1) = the slope the question provides all the info you need.

OpenStudy (anonymous):

ok

OpenStudy (trojanpoem):

If you don't understand don't feel ashamed, ask.